A student takes a ten-question true-false quiz, but did not study and randomly guesses each answer. Find the probability that the student passes the quiz with a grade of at least 40% of the questions correct. (Round your answer to three decimal places.)
Answer #1

Solution:- Given that n = 10, p = 1/2 = 0.5, q = 0.5 X = 10*0.4 = 4

P(X >= 4) = 1 - binomcdf(10,0.5,3) = 0.8281